# Universal Fine-Grained Symmetry Inference and Enforcement for Rigorous Crystal Structure Prediction

### v1

@@ A.4 Diffusion Rectification via Predicted Symmetry
$\hat{\mathbf{k}}_{t}=\mathbf{m}_{S_{g}}\odot\mathbf{k}_{t}+(\mathbf{1}-\mathbf{m}_{S_{g}})\odot\mathbf{k}_{ideal},$ (14)
where $\mathbf{k}_{ideal}$ represents the exact theoretical values required by the crystal family of $S_{g}$ (e.g., ensuring 90 angles or equal edge lengths).
Coordinate Rectification via Subspace Projection. In the absence of hard constraints, the fractional coordinates $\mathbf{F}_{t}\in\mathbb{R}^{N\times 3}$ generated at step $t$ inevitably violate the strict site symmetries derived in Step 2. We employ a Project-and-Reconstruct mechanism to rectify $\mathbf{F}_{t}$ into a symmetry-compliant state $\hat{\mathbf{F}}_{t}$.
@@
By simultaneously replacing the noisy states $\mathbf{k}_{t}$ and $\mathbf{F}_{t}$ with their rectified counterparts $\hat{\mathbf{k}}_{t}$ and $\hat{\mathbf{F}}_{t}$ at the end of each denoising step, we ensure that both the atomic configuration and the unit cell geometry strictly evolve on the geometric manifold defined by the target Wyckoff symmetries.
@@ Appendix B Hyper-parameters and Training/Inference Configurations
Complementing the mathematical formulations and algorithmic designs detailed in Supplementary Information A, which define the structural logic of our method, this part provides the granular implementation specifications necessary for reproducibility. We herein explicitly list the concrete hyperparameter valuations, optimization protocols, and inference settings used in our experiments. These configurations instantiate the abstract models into the specific computational pipelines employed to yield the reported results. Please note that the specifications listed in this section exclusively pertain to our proposed symmetry prediction method. Regarding the generative diffusion backbone, we strictly adhere to the original configurations of DiffCSP++ [13]. We utilized its official open-source implementation with identical hyperparameters, training schedules, and inference protocols, with the sole intervention being the substitution of the CSPML-retrieved templates with our ab initio predicted templates.
B.1 Data Representations and Tensor Shapes
To clarify the data flow described in the methodology, we formally define the shapes of the input and output tensors used in our implementation. The batch size $B$ is set as $32$. We set the maximum sequence length $L=128$ and maximum multiplicity slots $M=30$.
  • Inputs: The atomic sequence input is represented as $\mathbf{x}\in\mathbb{N}^{B\times L}$, with an associated padding mask $\mathbf{m}\in\{0,1\}^{B\times L}$. The conditioning space group labels are denoted as $\mathbf{s}\in\mathbb{N}^{B}$. For multiplicity constraints, we utilize a multiplicity sequence $\mathbf{u}\in\mathbb{N}^{B\times M}$ with a mask $\mathbf{m}_{u}\in\{0,1\}^{B\times M}$.
  • Embeddings: The output of the token embedding layer (Eq. 4 in Appendix A) corresponds to a tensor $\mathbf{E}(\mathbf{x})\in\mathbb{R}^{B\times L\times D}$, where $D=256$ is the hidden dimension.
  • Outputs: The model produces unnormalized logits $\mathbf{O}\in\mathbb{R}^{B\times L\times C}$, where $C=25$ is the size of the Wyckoff letter vocabulary. These logits form the basis for the probability distribution $p_{w}$ used in the beam search.
B.2 Model Architecture and MoE Settings
We employ a BERT-style Transformer encoder enhanced with Soft Mixture-of-Experts (SoftMoE) layers. The specific architectural hyperparameters are summarized in Table 3. Consistent with the notation in Appendix A, $D$ denotes the hidden embedding dimension, and $H$ represents the number of attention heads.
Table 3: Hyperparameter settings for the model architecture and SoftMoE layers.
@@ B.2 Model Architecture and MoE Settings
Space Group Tokens ($S_{g}$) 169 (include <unk>)
Wyckoff Labels ($C$) 25 (include <pad>)
B.3 Training and Testing Settings
The optimization protocols for training and the inference configurations for testing are unified in Table 4. We utilize the AdamW optimizer with a cosine annealing scheduler for training, and a constrained beam search with a fixed beam width for testing.
Table 4: Combined settings for optimization, training loop, and testing inference.
Category Hyperparameter Value
Optimization Optimizer AdamW
@@ B.3 Training and Testing Settings
Scheduling Scheduler ReduceLROnPlateau (monitor: val_loss)
@@
Testing & Inference Beam Width (K) 100
Search Strategy Constrained Beam Search
Sorting Criterion Log-likelihood (descending)
@@ C.1 Evaluation Metrics: The SUN and Matching Rate Metrics
Following the protocol defined by Zeni et al. [30], we evaluate generative quality using the SUN metrics. Leveraging the low inference cost of our model, we adopt a Top-20 evaluation strategy: for each test composition, we generate 20 candidate structures. Each candidate is assessed against three criteria:
  • Stability. A structure is classified as stable if its formation energy is within 0.1 eV/atom of the convex hull of competing phases. To ensure computational feasibility for large-scale generation, we follow Zeni et al. [30] and utilize MatterSim [27], a machine learning force field trained on extensive DFT trajectories, for geometry optimization and energy evaluation.
@@
Complementary to the SUN protocol, we employ the Matching Rate to evaluate the model’s ability to recover specific ground-truth (GT) structures. For each composition in the test set, we generate $k=20$ samples and identify a successful match if at least one sample is structurally equivalent to the corresponding GT structure. This structural equivalence is determined using the StructureMatcher class in pymatgen with a rigorous set of thresholds: fractional length tolerance $\text{stol}=0.5$, angle tolerance $\text{angle\_tol}=10^{\circ}$, and lattice vector tolerance $\text{ltol}=0.3$.
Accordingly, the Top-20 Matching Rate is defined as the proportion of matched compositions over the entire test set:
$\text{Top-20 Matching Rate}=\frac{N_{\text{matched}}@\text{Top-20}}{N_{\text{total}}}\times 100\%.$ (18)

### v2

@@ A.4 Diffusion Rectification via Predicted Symmetry
$\hat{\mathbf{k}}_{t}=\mathbf{m}_{S_{g}}\odot\mathbf{k}_{t}+(\mathbf{1}-\mathbf{m}_{S_{g}})\odot\mathbf{k}_{constrains},$ (14)
where $\mathbf{k}_{constrains}$ represents the exact theoretical values required by the crystal family of $S_{g}$ (e.g., ensuring 90 angles or equal edge lengths).
Coordinate Rectification via Subspace Projection. In the absence of hard constraints, the fractional coordinates $\mathbf{F}_{t}\in\mathbb{R}^{N\times 3}$ generated at step $t$ inevitably violate the strict site symmetries. We employ a Project-and-Reconstruct mechanism to rectify $\mathbf{F}_{t}$ into a symmetry-compliant state $\hat{\mathbf{F}}_{t}$.
@@
By simultaneously replacing the noisy states $\mathbf{k}_{t}$ and $\mathbf{F}_{t}$ with their rectified counterparts $\hat{\mathbf{k}}_{t}$ and $\hat{\mathbf{F}}_{t}$ at the end of each denoising step, we ensure that both the atomic configuration and the unit cell geometry strictly evolve on the geometric manifold defined by the target Wyckoff positions.
@@ Appendix B Hyper-parameters and Training/Inference Configurations
Gemini said This section details the implementation specifications for our proposed symmetry prediction method. For the generative diffusion backbone, we strictly adhere to the original configurations of DiffCSP++ [13]. We utilize its official open-source implementation with identical hyper-parameters, training schedules, and inference protocols, while replacing its CSPML-retrieved templates with our ab initio predicted Wyckoff templates to constrain its diffusion process. Table 3 summarizes the hyper-parameters for the model architecture, and Table 4 outlines the training and inference configurations.
Table 3: Hyper-parameters for the model architecture.
@@
Data & Tensor Shapes Batch Size ($B$) 32
Max Sequence Length ($L$) 128
Hidden Dimension ($D$) 256
Wyckoff Labels ($C$) 25 (include <pad>)
@@
Space Group Tokens ($S_{g}$) 231 (include <unk>)
Table 4: Training and inference configurations.
Hyperparameter Value
Optimizer AdamW
@@
Scheduler ReduceLROnPlateau
@@
Beam Width (K) 100
@@ C.1 Evaluation Metrics: The SUN and Matching Rate Metrics
Following the core principles of the protocol defined by Zeni et al. [30], we evaluate generative quality using the SUN metrics. Leveraging the low inference cost of our model, we adopt a Top-20 evaluation strategy: for each test composition, we generate 20 candidate structures. Each candidate is assessed against three criteria:
  • Stability. A structure is classified as stable if its formation energy is within 0.1 eV/atom of the convex hull of competing phases. While Zeni et al. [30] employed density functional theory (DFT) for the final stability verification of 1,024 samples, performing DFT relaxations across the entire test set is computationally prohibitive under our extensive Top-20 evaluation setting. To ensure computational feasibility, we utilize MatterSim [28], an accurate machine learning force field method validated by Zeni et al. [30], as a high-fidelity surrogate for geometry optimization and energy evaluation.
@@
Complementary to the SUN metrics, we employ the Top-20 Matching Rate to evaluate the model’s ability to recover ground-truth structures. For each testing composition, we generate 20 candidates, recording a successful match if at least one candidate is structurally equivalent to the GT. Exactly following previous works [12, 13], this equivalence is evaluated via the StructureMatcher class in pymatgen using tolerances of $\text{stol}=0.5$, $\text{angle\_tol}=10^{\circ}$, and $\text{ltol}=0.3$. The metric is then calculated as the percentage of matched compositions across the entire testing set:
$\text{Top-20 Matching Rate}=\frac{N_{\text{matched}}@\text{Top-20}}{N_{\text{total}}}\times 100$ (18)